OpenStudy (anonymous):

Question 1 The slope of the line containing the points (-5, 3) and (-2, 1) is? Question 2 The midpoint of A (-4, 2) and B(8, 5) is? Question 3 Indicate the equation of the line that is the perpendicular bisector of the segment with endpoints (4, 1) and (2, -5). Question 4 If the equation of a circle is (x + 5)2 + (y - 7)2 = 36, its radius is? Question 5 A trapezoid has symmetry with respect to the median of the trapezoid. true/false Question 6 The slope of a line that is perpendicular to a line whose slope is 3/8 is?

OpenStudy (anonymous):

N64, you should post one question at a time. Don't you think 6 questions at once is asking a bit too much of the community here?

OpenStudy (anonymous):

1. Use y2 - y1/ x2-x1 for the slope. 2. Use ((x1 + x2)/2, y1+y2/2)) 4. In a circle equation, (x-h)^2 +(y-k)^2 =r^2 6. To find a perpendicular line to another, you tape the reciprocal of the original slope.

OpenStudy (anonymous):

@Abs.Zero, I very much like the format of your answer. Good job.

OpenStudy (anonymous):

Questions 13 and 14 needs a diagram which is attach to it. Question 7 If the equation of a circle is (x - 2)2 + (y - 6)2 = 4, the center is point (2, 6). True/False Question 8 The equation of the line through point (- 4, 0) and parallel to the y-axis is? Question 9 A scalene triangle has symmetry with respect to a line containing an altitude of the triangle. True/False Question 10 If the equation of a circle is (x + 5)2 + (y - 7)2 = 36, its center point is? Question 11 Indicate the equation of the line through (2, -4) and having slope of 3/5. Question 12 Any line with no slope is parallel to the? Question 13 Tell if the slope is positive, negative, zero, or no slope. Question 14 In the figure, M and N are midpoints of RT and ST. What is the length of MN? Question 15 The equation of circle having a diameter with endpoints (-3, 1) and (5, 7) is? Question 16 Point (-4, 3) lies in Quadrant? Question 17 A trapezoid has all three types of symmetry. True/False Question 18 The distance between A (-5, 3) and B(5, 3) is?. Question 20 Point (6, 0) lies on the? Question 21 A square has symmetry with respect to a line containing a median of the square. TrueFalse Question 22 A parallelogram has symmetry with respect to the line that bisects it vertically. True/False Question 23 If the point (a,3) lies on the graph of the equation 5x + y = 8, then a=?

OpenStudy (anonymous):

@N64, we are not here to do your work for you.

OpenStudy (anonymous):

qns7. true. qns8. x=-4 qns9. true qns10. (-5,7) qns.11 -4 = (3/5) 2 +c c= -26/5 y= (3/5)x - 26/5 qns.12 y-axis qns.13 undefined qns.14 (2b-2a)/2 = b-a qns.15 circle has radius of 4, center point is (1,5) eqn of circle = (x-1)^2 +(y-5)^2 = 4^2 qns.16 quadrant 4 qns.17 false qns.18 10 qns.20 it lies on the x-axis qns.21 true qns.22 true qns23. 5x+y=8 > substitute (a,3) > 5a +3=8 > 5a = 5> a=1